Comfort Step Company manufactures and sells shoes. Comfort Step uses activity-based costing to determine the cost of the sales order processing and the shipping activity. The sales order processing activity has an activity rate of $72 per sales order, and the shipping activity has an activity rate of $20 per shipment. Comfort Step sold 10,000 units of walking shoes, which consisted of 5,400 orders and 500 shipments. a. Determine the total sales order processing and shipping activity cost for walking shoes. $
Answer #1

Total sales order processing cost = Sales orders * Rate per sales order =5400*72 = 388800

Total shipping activity cost = Shipments * Rate per shipment =500*20 =10000

Per unit sales order processing cost =Total cost / Units sold= 388800/10000 =38.88

Per unit shipping activity cost =10000/10000 =1